OpenStudy (anonymous):

Which equation is a quadratic equation? y + 3 = 5x + 10 y + 4x2 = (4x − 2)(x + 3) y − 8x = (x2 + 1)(9x − 7) y = 2x + 7x(x + 1)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@phi

OpenStudy (phi):

what makes an equation "quadratic" ?

OpenStudy (anonymous):

The squared variable?

OpenStudy (phi):

its highest degree (exponent) is 2. It has a 2nd degree term squared variable YES

OpenStudy (phi):

so if there is only x^1 or no x, cross it off the list.

OpenStudy (anonymous):

That leaves the middle two.

OpenStudy (phi):

if you multiply x times x^2 or higher you will get x^3 or higher. that is too high. cross those off the list

OpenStudy (phi):

I would not cross off the bottom multiplying x*x gives x^2 so that is ok

OpenStudy (anonymous):

The third one?

OpenStudy (phi):

the 3rd one has (x^2 + 1)(9x − 7) notice you will end up multiplying x^2 times 9x to get 9x^3 no. the exponent is 3 (too big)

OpenStudy (phi):

it will be 2nd or 4th the 4th is definitely quadratic... you can multiply it out and simplify and it looks like a very respectable quadratic (you could take it home to meet the folks)

OpenStudy (anonymous):

LOL, love the last line. Thanks!

OpenStudy (phi):

the 2nd one is tricky. can you multiply it out and simplify it ? y + 4x^2 = (4x − 2)(x + 3)

OpenStudy (anonymous):

I forgot how to do that.

OpenStudy (phi):

on the right side of the =, use FOIL (4x -2)(x+3) First: 4x * x = 4x^2 Outer: 4x * 3= 12x Inner: -2 * x = -2x Last: -2*3 = -6 add up: 4x^2 + 10 x -6 so you now have y + 4x^2 = 4x^2 + 10 x -6 can you simplify to y = .... ?

OpenStudy (phi):

you would add -4x^2 to both sides \[ y + 4x^2 -4x^2 = 4x^2 -4x^2 + 10 x -6 \] can you simplify ?